OpenStudy (anonymous):

• What are the differences among expressions, equations, and functions? Provide examples of each.

OpenStudy (anonymous):

I just to not get math can anyone help me please

OpenStudy (anonymous):

In mathematics, an expression is a general term for any well formed combination of mathematical symbols. It may consist of just one or more numbers, variables, operators, functions, and their arguments that evaluate to a single answer. [ Equations define the relationship between one or more variables. Single variable equation: x2 - 10x = 0 Multi vairable equation: 2x+3y = 10 An equation doesnt output a value. We can solve the equation and find out the values of the variables. The equation would be valid only for certain values of the variables. A function always outputs a value. Suppose the function f(x) = 10x+20, what this means is that this function outputs different values for different input values of variable 'x'. For example: f(10) = 10X10+20 = 120 f(20) = 10X20+20 = 220 i.e. we substitute different values into x and get different outputs. The important point is that every function outputs a value ]